# MINDSTORES: Memory-Informed Neural Decision Synthesis for Task-Oriented Reinforcement in Embodied Systems

### v1

@@ 7 Conclusion
Impact Statement
This work introduces a novel approach to autonomous Minecraft gameplay by combining large language models with dynamic experience storage. This system demonstrates human-like problem-solving capabilities in complex and safe open-world environments by breaking down high-level goals into executable actions through natural language reasoning. The architecture’s ability to learn from past experiences and adapt to new scenarios represents a significant step toward more versatile and intelligent game-playing agents. This research is broadly applicable beyond gaming to other domains including real-world robotic tasks and autonomous systems, where additional consideration of safety developments would be needed to create physical systems.
A Method
@@ A.1 Agent Algorithm
Pseudocode 1: Agent algorithm.
def run_agent(
    environment,    # MineDojo environment
    max_steps=1000, # Maximum steps to run
    goal_input=""   # Optional high-level goal
    # Initialize metrics and experience tracking
    metrics_logger = MetricsLogger()
    experience_store = ExperienceStore()
    # Initial environment reset
    obs, _, _, info = environment.step(environment.action_space.no_op())
    step = 0
    while step < max_steps:
        # 1. Create structured state description
        state_json = get_state_description(obs, info)
        # 2. Get next immediate task
        sub_task = get_next_immediate_task(state_json)
        metrics_logger.start_subtask()
        # 3. Plan action sequence
        actions = plan_action(state_json, info["inventory"], sub_task)
        # 4. Execute actions and track experience
        obs, reward, done, info = execute_action_sequence(actions)
        # 5. Store experience and update metrics
        if done:
            store_experience(state_json, reward, done)
            break
        step += len(actions)
    environment.close()
    metrics_logger.print_summary()
@@ A.2 LLM Prompts
*A.2.1 Environment Description Prompt
You are an expert Minecraft observer. Describe the current environment state focusing on:
1. The agent’s immediate surroundings (blocks, entities, tools)
2. Environmental conditions (weather, light, temperature)
3. Agent’s physical state (health, food, equipment)
4. Notable resources or dangers
Current state:
\${state_json_str}
Provide a clear, concise description that would be useful for planning actions.
*A.2.2 Situation Analysis Prompt
You are an expert Minecraft strategist. Given the current state and environment description:
1. Analyze available resources and their potential uses
2. Identify immediate opportunities or threats
3. Consider crafting possibilities based on inventory
4. Evaluate progress towards goals
Environment description:
\${description}
Current state:
\${state_json_str}
Provide strategic insights about the current situation.
*A.2.3 Strategy Planning Prompt
You are an expert Minecraft planner. Create a strategic plan considering:
1. The current goal: \${goal}
2. Available resources and tools
3. Environmental conditions
4. Potential obstacles or requirements
5. Do not assume intermediate tasks can be achieved without running another agent loop
6. Specify quantities and required actions
Environment description:
\${description}
Situation analysis:
\${explanation}
Current state:
\${state_json_str}
Create a specific, actionable plan that moves towards the goal.
*A.2.4 Action Selection Prompt
You are an expert Minecraft action selector. Convert the plan into specific actions:
1. Use only valid Minecraft actions (move_forward, move_backward, jump, craft, etc.)
2. Consider the current state and available resources
3. Break down complex tasks into simple action sequences
4. Ensure actions are feasible given agent capabilities
5. Make actions incremental and build progressively
Available actions:
- forward [N]: Move forward N steps (default 1)
- backward [N]: Move backward N steps (default 1)
- move_left
- move_right
- jump
- sneak
- sprint
- attack [N]
- use
- drop
- craft
- equip [item]
- place [block]
- destroy
- look_horizontal +/-X
- no_op
Strategic plan:
\${plan}
Current state:
\${state_json_str}
Return ONLY a list of actions, one per line, that can be directly executed.
*A.2.5 Outcome Evaluation Prompt
Evaluate the outcome of a Minecraft action sequence in brief.
Initial state (JSON): \${initial_state}
Final state (JSON): \${final_state}
Reward: \${reward}
Done: \${done}
GPT Plan: \${gpt_plan}
Executed Actions: \${executed_actions}
Format response as: outcome|success|explanation
A.3 State Representation
The structured state representation includes:
Core Components:
  • Inventory: Dictionary mapping items to quantities
  • Equipment: Currently equipped armor/weapons/tools
  • Nearby blocks: Block types within 32-block radius
  • Position: 3D coordinates in world space
  • Health/Hunger bars: Current status (max 20)
Environmental Information:
  • Biome type and characteristics
  • Time of day (sunrise, day, noon, sunset, night, midnight)
  • Weather conditions
  • Light levels
  • Nearby entities with distances
Task Tracking:
  • Completed tasks history
  • Failed tasks log
  • Current active subtask
  • Task dependencies
A.4 Experience Store
The experience store maintains a database of past experiences with the following structure:
@dataclass
class ActionExperience:
    state: str           # Initial state description
    task: str           # Attempted task
    plan: List[str]     # Action sequence
    outcome: str        # Result description
    success: bool       # Task completion status
    reward: float       # Numerical reward
    embedding: np.ndarray # State embedding vector
Key functionality includes:
  • Semantic search using SBERT embeddings
  • Experience retrieval based on state/task similarity
  • Automatic logging of outcomes
  • Database health monitoring
  • Experience pruning based on relevance
A.5 Metrics Logger
The metrics logger tracks:
  • Total subtasks attempted
  • Successful subtasks completed
  • Experience retrieval statistics
  • Subtask completion times
  • Database size over time
  • Action success rates
  • Resource collection efficiency
Example metrics output:
===== METRICS SUMMARY =====
Total subtasks: 47
Successful subtasks: 32
Success rate: 68.1%
Experience retrieval calls: 94
Total experiences retrieved: 283
Avg subtask completion time: 12.4s
Database size: 156 experiences
Action success rate: 73.2%
Resource efficiency: 82.5%
B Implementation Details
@@ B.1 Core Components
The codebase is structured into core modules:
  • State representation and processing
  • Experience management and retrieval
  • Action planning and execution
  • Metrics collection and analysis
  • Environment interaction handlers
@@ B.2 Environment Integration
Environment configuration:
env = minedojo.make(
    task_id="survival",
    image_size=(480, 768),
    seed=40,
    initial_inventory=[
        InventoryItem(slot=0, name="wooden_axe", quantity=1),
Action space includes:
  • Movement: forward, backward, left, right, jump, sneak, sprint
  • Interaction: attack, use, drop, craft, equip, place, destroy
  • Camera: look_horizontal, look_vertical
  • Special: no_op
@@ B.3 Neural Components
  • Model: SBERT ’all-MiniLM-L6-v2’
  • Output dimension: 384
@@
  • Distance metric: Euclidean
@@
C Appendix
C.1 Configuration Parameters
DEFAULT_CONFIG = {
    "max_steps": 1000,
    "max_recent_actions": 5,
    "experience_retrieval_k": 5,
    "embedding_dim": 384,
    "max_retries": 4,
    "action_timeout": 300,
    "exploration_rate": 0.1
C.2 Logging Format
logging.basicConfig(
    format=’%(asctime)s - %(name)s - %(levelname)s - %(message)s’,
    level=logging.INFO,
    handlers=[
        logging.FileHandler("agent.log"),
        logging.StreamHandler(sys.stdout)
C.3 Error Handling
Exception hierarchy:
class AgentError(Exception): pass
class PlanningError(AgentError): pass
class ExecutionError(AgentError): pass
class ExperienceError(AgentError): pass
Recovery strategies:
  • Automatic retry with exponential backoff
  • Fallback to simpler actions
  • State restoration on critical failures
  • Experience logging for failed attempts
@@ C.3 Error Handling
Table 2: Task Details with MINDSTORES and DEPS Percentages
k Torch Iron Boots Iron Pickaxe Minecart Diamond Context Description Avg Time
1 3.3 10 0 0 0 Minimal context 3 s
3 6.6 20 3.3 0 0 Slight overhead 15 s
5 13.3 27 6.67 3.3 0 Sweet spot 23 s
10 23.3 33.33 10 6.67 0 Overcrowded 57 s
20 23.3 33.3 13.3 6.67 0 Leveled off 110 s
Table 3: Retrieved Experiences for Different Values of k
Task Number Task Name Novel DB Size Success Steps
1 Wooden Door 26 Yes 3000
2 Wooden Shovel 56 Yes 4357
3 Furnace 81 Yes 4879
4 Cooked Beef 134 Yes 5602
5 Cooked Porkchop 141 Yes 5802
6 Torch 197 Yes 6458
7 Stone Slab 199 Yes 6578
8 Iron Pickaxe 289 Yes 8598
9 Iron Axe 300 Yes 8986
10 Minecart 355 Yes 9112
Table 4: Task Details with Database Size, Success, and Steps till Completion
@@ 10 Minecart 355 Yes 9112
Make Stone Sword 187 263 N/A (500+)
Mine Iron 276 N/A (500+) N/A (500+)
Table 5: Time steps required to complete different Minecraft tasks across three systems
@@
Table 6: Success Rate Comparison With vs. Without Outcome Prediction (MINDSTORES vs. DEPS)
Note: Data corresponds to the performance comparison graph. MINDSTORES shows consistent improvements across all tasks (mean +9.4%). Performance gap widens with complexity until MT8 (terminal failure for both).

### v2

@@ 7 Conclusion
Appendix A: Agent Algorithm and LLM Prompts
@@ A.1 Agent Algorithm
    1 def run_agent( 2     environment,    # MineDojo environment 3     max_steps=1000, # Maximum steps to run 4     goal_input=""   # Optional high-level goal 5 ): 6     # Initialize metrics and experience tracking 7     metrics_logger = MetricsLogger() 8     experience_store = ExperienceStore() 9  10     # Initial environment reset 11     obs, _, _, info = environment.step(environment.action_space.no_op()) 12  13     step = 0 14     while step < max_steps: 15         # 1. Create structured state description 16         state_json = get_state_description(obs, info) 17  18         # 2. Get next immediate task 19         sub_task = get_next_immediate_task(state_json) 20         metrics_logger.start_subtask() 21  22         # 3. Plan action sequence 23         actions = plan_action(state_json, info["inventory"], sub_task) 24  25         # 4. Execute actions and track experience 26         obs, reward, done, info = execute_action_sequence(actions) 27  28         # 5. Store experience and update metrics 29         if done: 30             store_experience(state_json, reward, done) 31             break 32  33         step += len(actions) 34  35     environment.close() 36     metrics_logger.print_summary()
@@ A.2 LLM Prompts
A.2.1 Environment Description Prompt
    1 You are an expert Minecraft observer. Describe the current environment state focusing on: 2 1. The agent’s immediate surroundings (blocks, entities, tools) 3 2. Environmental conditions (weather, light, temperature) 4 3. Agent’s physical state (health, food, equipment) 5 4. Notable resources or dangers 6  7 Current state: 8 \${state_json_str} 9  10 Provide a clear, concise description that would be useful for planning actions.
A.2.2 Situation Analysis Prompt
    1 You are an expert Minecraft strategist. Given the current state and environment description: 2 1. Analyze available resources and their potential uses 3 2. Identify immediate opportunities or threats 4 3. Consider crafting possibilities based on inventory 5 4. Evaluate progress towards goals 6  7 Environment description: 8 \${description} 9  10 Current state: 11 \${state_json_str} 12  13 Provide strategic insights about the current situation.
A.2.3 Strategy Planning Prompt
    1 You are an expert Minecraft planner. Create a strategic plan considering: 2 1. The current goal: \${goal} 3 2. Available resources and tools 4 3. Environmental conditions 5 4. Potential obstacles or requirements 6 5. Do not assume intermediate tasks can be achieved without running another agent loop 7 6. Specify quantities and required actions 8  9 Environment description: 10 \${description} 11  12 Situation analysis: 13 \${explanation} 14  15 Current state: 16 \${state_json_str} 17  18 Create a specific, actionable plan that moves towards the goal.
A.2.4 Action Selection Prompt
    1 You are an expert Minecraft action selector. Convert the plan into specific actions: 2 1. Use only valid Minecraft actions (move_forward, move_backward, jump, craft, etc.) 3 2. Consider the current state and available resources 4 3. Break down complex tasks into simple action sequences 5 4. Ensure actions are feasible given agent capabilities 6 5. Make actions incremental and build progressively 7  8 Available actions: 9 - forward [N]: Move forward N steps (default 1) 10 - backward [N]: Move backward N steps (default 1) 11 - move_left 12 - move_right 13 - jump 14 - sneak 15 - sprint 16 - attack [N] 17 - use 18 - drop 19 - craft 20 - equip [item] 21 - place [block] 22 - destroy 23 - look_horizontal +/-X 24 - no_op 25  26 Strategic plan: 27 \${plan} 28  29 Current state: 30 \${state_json_str} 31  32 Return ONLY a list of actions, one per line, that can be directly executed.
A.2.5 Outcome Evaluation Prompt
    1 Evaluate the outcome of a Minecraft action sequence in brief. 2  3 Initial state (JSON): \${initial_state} 4 Final state (JSON): \${final_state} 5 Reward: \${reward} 6 Done: \${done} 7 GPT Plan: \${gpt_plan} 8 Executed Actions: \${executed_actions} 9  10 Format response as: outcome|success|explanation
Appendix B: Implementation Details
@@ B.1 Core Components
The codebase is structured into modules for state processing, experience management, action planning, metrics collection, and environment interaction.
@@ B.2 Environment Integration
    1 env = minedojo.make( 2     task_id="survival", 3     image_size=(480, 768), 4     seed=40, 5     initial_inventory=[ 6         InventoryItem(slot=0, name="wooden_axe", quantity=1), 7     ] 8 )     The action space includes movement (forward, backward, left, right, jump, sneak, sprint), interaction (attack, use, drop, craft, equip, place, destroy), camera control (look_horizontal, look_vertical), and special (no_op).
@@ B.3 Neural Components
  • Model: SBERT ‘all-MiniLM-L6-v2’
  • Output dimension: 768
@@
  • Distance metric: cosine similarity
@@
Appendix C: Additional Tables
@@
Table 2: Task Details with MINDSTORES and DEPS Percentages
@@
Make Stone Sword 187 263 500
Mine Iron 276 500 500
Table 3: Time steps required to complete different Minecraft tasks across three systems. (Values for Voyager and Reflexion are capped at 500 in some tasks.)
@@
Table 4: Success rate comparison with outcome prediction (MINDSTORES) vs. without (DEPS).